OpenStudy (anonymous):

Hi I'm new to open study, If someone could please help me with these 3 questions. I cant seem to get them right. 1) Given the following functions f(x) and g(x), solve for (f ⋅ g)(2) : f(x) = 3x2 + 2 g(x) = x − 8 2) Given the following functions f(x) and g(x), solve f[g(10)] f(x) = 10x + 8 g(x) = x + 9 3) Given the following functions f(x) and g(x), solve (f + g)(3) f(x) = 2x + 21 g(x) = x − 24

OpenStudy (phi):

(f ⋅ g)(2) is short for f(2) * g(2) can you find f(2) ?

OpenStudy (anonymous):

I have no idea how to start.

OpenStudy (phi):

they give you a formula f(x) = 3x^2 + 2 f(x) is the "short name" and 3x^2 + 2 is the formula f(2) means: find the formula, find the "x" in the formula and replace the x with a 2 can you do that?

OpenStudy (anonymous):

So 3(2)^2+2?

OpenStudy (anonymous):

38?

OpenStudy (phi):

yes. now simplify that

OpenStudy (anonymous):

Is it 38?

OpenStudy (phi):

no. how did you get that? remember order of operations first do exponents, so first do 2^2

OpenStudy (phi):

2^2 means 2*2

OpenStudy (anonymous):

Well I replaced the x with 2. And the formula is 3x^2 + 2. Is it 16?

OpenStudy (anonymous):

3 times 4 is 12 and 12 plus 2 is 14. Sorry 14?

OpenStudy (phi):

it is worth knowing how to do this. 3*2^2 +2 first do 2^2 which is 2*2 . you get 4 so you have 3*4 +2 12+2 14 yes 14

OpenStudy (phi):

(f ⋅ g)(2) is short for f(2) * g(2) so far we have f(2) is 14 now we need g(2) can you do that ?

OpenStudy (anonymous):

Where do I start?

OpenStudy (phi):

to find g(2) find the formula for g replace any x's in the formula with 2 simplify

OpenStudy (anonymous):

How do I find the formula for g?

OpenStudy (anonymous):

Oh they gave me the formula

OpenStudy (anonymous):

g(x) = x − 8

OpenStudy (phi):

they give it to you. g(x) is the "name" they give to x-8 that is what g(x) = x − 8 means

OpenStudy (anonymous):

So g(2)=2-8

OpenStudy (phi):

yes, and that simplifies to what?

OpenStudy (anonymous):

g(2)=-6

OpenStudy (phi):

ok now we have (f ⋅ g)(2) is short for f(2) * g(2) so far we have f(2) is 14 and g(2) is -6 now we multiply to get the final answer

OpenStudy (anonymous):

14*-6=-84

OpenStudy (phi):

yes.

OpenStudy (anonymous):

Thankyou! Can you help with the other 2?

OpenStudy (phi):

Let's skip to Q3 (f + g)(3) f(x) = 2x + 21 g(x) = x − 24 can you guess what (f+g)(3) is short for ? (based on the first question)

OpenStudy (anonymous):

f(3) and g(3)

OpenStudy (phi):

it would be better to say f(3) + g(3)

OpenStudy (anonymous):

f(3)=2(3)+21 f(3)=27?

OpenStudy (phi):

yes

OpenStudy (anonymous):

So g(x) = x − 24 g(3)=3-24 g(3)= -21?

OpenStudy (phi):

yes

OpenStudy (anonymous):

And 27*-21 is -567

OpenStudy (phi):

ok, but we are doing (f + g)(3) = f(3) + g(3)

OpenStudy (anonymous):

Ok so what do we do know that we know what f and g are?

OpenStudy (phi):

the + means add

OpenStudy (anonymous):

27+(-21)=6

OpenStudy (phi):

yes. notice in Q1 we have multiply (that is what the dot means) and in this question we are adding we might also see (f-g)(3) (subtract) or (f/g)(3) divide

OpenStudy (phi):

now Q2 this one is a bit different. The formulas are not combined using add or multiply we have something called "composition" (they have to call it something)

OpenStudy (anonymous):

Im still stuck on 3

OpenStudy (anonymous):

My answer choices are: 98 198 190 2052

OpenStudy (phi):

those are the choices for (f + g)(3) where f(x) = 2x + 21 g(x) = x − 24 ??

OpenStudy (anonymous):

Yes

OpenStudy (phi):

those look like the choices for Q2 which we have not done yet

OpenStudy (anonymous):

Oh Im sorry I got confused

OpenStudy (anonymous):

The answer was 6 sorry

OpenStudy (anonymous):

Ok we can do #3 now

OpenStudy (phi):

Q2 solve f[g(10)] f(x) = 10x + 8 g(x) = x + 9 think of g(10) as some number (which it is) let's find g(10) can you do that ?

OpenStudy (anonymous):

Ok so g(10)=10+9?

OpenStudy (phi):

yes, and that simplifies to 19 so g(10) is 19 if we replace g(10) with 19 in f[g(10)] we get f(19) (the [ vs ( is just to make it easier to read in f[g(10)] ... but it means the same as parens) anyway we have figured out that f[g(10)] is the same as f(19) now we need to find f(19)

OpenStudy (anonymous):

f(x) = 10x + 8 So f(19)=10(19)+8?

OpenStudy (phi):

ok, and what number do you get?

OpenStudy (phi):

you do 10 times 19 then add 8

OpenStudy (anonymous):

198?

OpenStudy (anonymous):

Thankyou so so much. I got a 100%! I havent gotten a 100% before. You're a really good tutor!